10a2b5c-5 (a to the 2nd power b to the 5th power and c to the -5th power)
__________ (over or divided by)
2a2b-4c4 (a to the 2nd power b to the -4th power and c to the 4th power)

Answer by checkley71(8403)

LETS TAKE THIS ONE ELEMENT AS A TIME:
10/2=5
A^2/A^2=1
B^5/B-4=B^(5+4)=B^9
C^-5/C^4=1/C^4+5)=1/C^9
5*1*B^9/C^9
5B^9/C^9 ANSWER.